OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. -2y^2 + y=-9

OpenStudy (anonymous):

The quadratic equation is \[x=(-b \pm \sqrt{b^2-4ac}\div2a\] Your equation can be rewritten in the form y=ax^2+bx+c which gives you the vales to sub in

OpenStudy (amistre64):

\frac{top}{bottom} might be useful

OpenStudy (amistre64):

\[\frac{\color{purple}{top}}{\color{green}{bottom}}\]tada

OpenStudy (anonymous):

uhh

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Thank you.

OpenStudy (amistre64):

\[\color{#af0c78}{\text{ Your welcome :)}}\]